Using 10 mL of 0.2 M H2SO4 solution, calculate the volume of water necessary to create a
0.05 M solution of H2SO4. What is the final volume of the solution?
1
Expert's answer
2020-05-08T14:11:34-0400

CM= n/V.

n=CM x V = 0.2 x 0.01 = 0.002

V=n/CM=0.002/0.05 = 0.04 L.

∆V=40 - 10 = 30 ml of necessary of water.